OpenStudy (anonymous):

How do you start simplifying the equation (cos(x)cot(x)/1-sin(x))-1?

OpenStudy (anonymous):

\[\frac{ \cos(x)\cot(x) }{ 1-\sin(x) }-1\] is this your question.

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

you can complete the answer.

OpenStudy (anonymous):

Thanks!

OpenStudy (loser66):

to me, I break the first term and the second term apart. second term is 1, let it there. now calculate the first term: \[\frac{cos *\frac{cos}{sin}}{1-sin}= \frac{\dfrac{cos^2}{sin}}{1-sin}\]time both numerator and denominator by (1+sin) \[\frac{\frac{cos^2}{sin}(1+sin)}{(1-sin)(1+sin)}=\frac{\frac{cos^2}{sin}(1+sin)}{1-sin^2}=\frac{\frac{cos^2}{sin}(1+sin)}{cos^2}\]cancel out cos^2 \[\frac{1+sin}{sin}=\frac{1}{sin}+1= csc +1\] now, add the second term from the beginning, that is -1 csc +1-1= csc